1500 mL of 0.9% NS in 12 hours. The administration set is 10 gtts/mL. Calculate the gtts/min. _1500 mL of 0.9% NS in 12 hours. The administration set is 10 gtts/mL. Calculate the gtts/min. _
Answer #1

1500 mL of 0.9% NS is administered in 12 hours.

The drop factor is at 10 gtts/mL

which means 10 drops per 1 ml.

for 1500 ml it is = 1500*10 gtts/12 hours

which is 15,000 gtts/12 hours

for 1 hour it is 15000/12 = 1250 gtts/ hr

for one minute it is = 1250/60 gtts /min

It is = 20.83 gtts/min which is 21 gtts/min approximately

This is a brief explanation. Hope this helps you out. Please feel free to give a like and contact us if you have further queries. Thank you and have a great day.